OpenStudy (anonymous):

Given the system of equations, what is the value of the system determinant? x + y = 8 x - y = 10

OpenStudy (anonymous):

To find the determinant of a 2x2 matrix: \[A = \left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\\ |A|=det(A)=ad-cb\]

OpenStudy (anonymous):

I'd actually forgotten how to do this since it's been so long, and it literally took a 2 second Google search ;)

OpenStudy (anonymous):

a and c is the value of x and b and d is the value of y

OpenStudy (anonymous):